# You Cannot Feed Two Birds with One Score: the Accuracy-Naturalness Tradeoff in Translation

### v1

@@ 3.1 The Tradeoff
To make the inevitability of the tradeoff more concrete, consider translating the German sentence “Meine Lehrerin ist nett” to English. A sensible translation would be “My teacher is nice,” which is fully natural, but loses the information from the German sentence that the teacher is female. An alternatively translation could be “My female teacher is nice,” which is now fully accurate, but sounds quite unnatural in English. Another, naturally occurring example of this phenomenon is the translation of Japanese honorifics. A completely accurate translation into English would need to introduce diverse paraphrases denoting the social relation between the speakers, which would result in unnatural text. This type of strict semantic differentiation is currently not explicitly assessed by any metric that we are aware of, but might be relevant in certain scenarios.
@@ 4 Translation evaluation in practice
Before moving to the empirical analysis, we lay out the theoretical basis for a practical evaluation pipeline. Thus, let ${\mathbf{X}}=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\ldots,{\mathbf{x}}_{N}\}$ and ${\mathbf{Y}}^{ref}=\{{\mathbf{y}}_{1},{\mathbf{y}}_{2},\ldots,{\mathbf{y}}_{N}\}$ be a parallel test set, with source-reference pairs drawn iid: ${\mathbf{x}}_{i},{\mathbf{y}}_{i}\sim P_{{\mathbf{x}},{\mathbf{y}}}$. Furthermore, let $R_{\mathbf{y}}$ be a monolingual reference distribution in the target language, which we could obtain by estimating it from some monolingual corpus, for example. Finally, we consider a set of translation systems to be evaluated and which we wish to compare with one another: $\mathcal{Q}=\{Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{1},\ldots Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{M}\}$. Following standard practice, we are interested in comparing models based on their outputs. Concretely, we require that each system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{m}\in\mathcal{Q}$ translates the source texts in the test set and produces a translation set ${\mathbf{Y}}^{m}=\{{\mathbf{y}}_{n}\sim Q_{{\mathbf{y}}\mid{\mathbf{x}}_{n}}^{m}\mid{\mathbf{x}}_{n}\in{\mathbf{X}}\}$. Our goal then, is to estimate the accuracy and naturalness of each of the $M$ systems using ${\mathbf{X}}$, ${\mathbf{Y}}^{ref}$, $R_{\mathbf{y}}$ and the ${\mathbf{Y}}^{m}$s.
@@ 4.1 No-reference metrics as statistical distances
So far, we followed a bottom-up approach, and developed our theory of accuracy and naturalness from the ground up. In this section, we switch to a top-down view, and ask whether current evaluation practices for assessing naturalness conform to our theory. The prevailing practical approach for data quality assessment are no-reference metrics. These can be either human-generated, such as the fluency scores derived from MQM assessments for text [^3] Strictly speaking MQM evaluation of fluency is not completely reference-free, as the judges have access to the source sentence. But fluency-related scores are judged mainly on a monolingual level. (Freitag et al., 2021), or automated, such as the perplexity of a language model for text. We formalise no-reference metrics by assuming we have a set of critics $\mathcal{C}=\{f_{1},f_{2},\ldots f_{K}\}$ (usually with $K=1$), each of which assigns a score to a test set ${\mathbf{Y}}$. Then, for each $f_{k}\in\mathcal{C}$, we compute the score $s_{k}=N^{-1}\sum_{n=1}^{N}f_{k}({\mathbf{y}}_{n})$ and average the result to get the final score $s=K^{-1}\sum_{k=1}^{K}s_{k}$. Within the context of MQM assessment, the critics could be the set of human raters evaluating the output of a translation system. Then, $f_{k}({\mathbf{y}}_{n})$ could be the $k$th rater’s fluency score for the $n$th hypothesis in the test set derived from their MQM evaluation according to the conversion scheme of Freitag et al. (2021). Does this evaluation procedure fit our definition on naturalness? If so, how? As an answer to this question we propose the following statistical distance between two distributions $Q$ and $R$:
$D_{1}(Q,R\mid\mathcal{P})=\mathbb{E}_{f\sim\mathcal{P}}\left[\left\lvert\mathbb{E}_{X\sim Q}[f(X)]-\mathbb{E}_{X\sim R}[f(X)]\right\rvert\right]$ (4)
where $\mathcal{P}$ is a probability distribution over critics (technically, it is a stochastic process). Note the similarity of Equation 4 to integral probability metrics (Müller, 1997): instead of maximising, we average over critics instead; see Appendix A for a detailed discussion. Formally defining $D_{1}$ is somewhat technical, hence we relegate the details of the definition and the analysis of basic properties to Appendix E. Here we argue that averaging no-reference scores is, in fact, a rough Monte Carlo estimate of an appropriately chosen $D_{1}$. Indeed, assuming that the critic scores are all non-negative and are such that the reference distribution $R_{\mathbf{y}}$ always achieves a perfect score for any $f\sim\mathcal{P}$, that is, $\mathbb{E}_{X\sim R_{\mathbf{y}}}[f(X)]=0$, then for a test set ${\mathbf{Y}}^{m}$ we have $D_{1}(Q^{m}_{\mathbf{y}},R_{\mathbf{y}}\mid\mathcal{P})\approx(KN)^{-1}\sum_{k=1}^{K}\sum_{{\mathbf{y}}\in{\mathbf{Y}}^{m}}^{N}f_{k}({\mathbf{y}})$ with $f_{k}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{P}$, which is precisely the score $s$ we obtain from averaging no-reference metrics.
@@ 5.1 Approximating the Accuracy-Naturalness curve using LLMs
First, we aim to get a basic feel for the accuracy-naturalness curve. Unfortunately, performing the optimisiation in Equation 3 is hopeless. Hence, we approximate it in the following way: using Gemini 1.5 Flash (Gemini Team et al., 2024), we generated $K=1024$ candidate translations $\{{\mathbf{y}}_{i}^{k}\}_{k=1}^{K}$ for each source sentence ${\mathbf{x}}_{i}$ in the WMT24 en $\to$ de test set; see Appendix F for the precise details. Next, we chose to two accuracy metrics $\Delta$ to test for: chrF (Popović, 2015) and Comet (Rei et al., 2020). For the naturalness metric, we used the pretrained Gemma2 2B model log-perplexities (Gemma Team et al., 2024). Concretely, let $\Gamma(y)$ denote the probability of a sequence of tokens $y$ as predicted by the Gemma2 model, let $\lvert y\rvert$ denote the number of tokens in $y$ and hence let $\mathrm{LPP}(Q)=\mathbb{E}_{Y\sim Q}\left[-\log\Gamma(Y)/\lvert Y\rvert\right]$ denote the log-perplexity of a distribution $Q$ with respect to $\Gamma$. Then, our concrete instantiation of $D_{1}$ is $D(Q,R)=\left\lvert\mathrm{LPP}(Q)-\mathrm{LPP}(R)\right\rvert$. As we show in Section E.1.2, this quantity can also be viewed as a more general version of the $D_{1}$ distance. We estimated the monolingual target reference distribution $R_{\mathbf{y}}$, using 20,000 entries of the 2024 NewsCrawl DEU Monolingual dataset [^5] Available at https://data.statmt.org/news-crawl/de/.. In our experiments, we found that $\mathrm{LPP}(R_{\mathbf{y}})$ was always smaller than $\mathrm{LPP}(Q^{m}_{\mathbf{y}})$, for each system we examined, hence $D$ is equivalent up to an additive constant to the average model log-perplexity on the test dataset $D(Q_{\mathbf{y}},R_{\mathbf{y}})=\mathrm{LPP}(Q_{\mathbf{y}})+\mathcal{O}(1)$. To optimise Equation 3, we first construct the objective’s one-shot Lagrange dual with tradeoff parameter $\beta$:
@@
We plot the results in Figure 1 and see that it verifies our theory: 1) here is indeed a tradeoff curve (see also Theorem B.1) and 2) the curve is non-increasing and concave, as 3.1 predicts. Finally, note that curves produced this way overestimates the true AN function and hence is overly optimistic, see Appendix F for the details.
@@ 5.2 Evaluating existing translation systems using the accuracy-naturalness framework
The plots of the previous section can be considered a simulation of the tradeoff curve through a controlled experiment, but how do real systems behave with respect to the ideal curve? In Figure 2, we include points for the systems that participated in the shared task in addition to the previously computed curve. We can see that all the systems are below the curve, as expected. Furthermore, the best performing systems in the evaluation (according to human judgement) are the ones that are closer to the curve. In order to investigate this assertion in more detail, in Figure 3 we show similar plots using human fluency and adequacy scores [^6] See Appendix D for details on the computation of fluency and adequacy scores from MQM evaluations. for those languages for which MQM evaluations are available. Note that in this case we cannot compute (or simulate) the tradeoff curve for human judgements.
@@
Figure 2: System performances on the WMT24 en $\to$ de test set. $\bullet$ marks represent LLMs, $\blacksquare$ marks represent systems specifically trained for MT, $\blacktriangle$ marks represent “online” systems and the $\blacklozenge$ represents the performance of an alternative human reference. For further explanation of these categories, please see Kocmi et al. (2024). We also include the approximate AN curves from Figure 1, though note that it is an optimistic estimate of the real curve.
Figure 3: Plot of the accuracy-naturalness tradeoff derived from human MQM adequacy and fluency ratings, as described in Section 5.2. We highlight the top-performing systems at the Pareto frontier.
@@
Secondly we can focus on interpreting the results for the performance of single systems. Unbabel (Rei et al., 2024) had the dominant system in the automatic metrics, but in human evaluations other systems were judged better. Unbabel’s system was clearly optimized for Comet, an accuracy measure. Looking at the plots in Figure 3 we can see that the Unbabel system has a quite strong adequacy score, but due to the tradeoff the fluency of the system suffers. Other systems that human raters judged to be better lie closer to the tradeoff curve. See Appendix G for comparisons of further automated metrics, as well as comparisons of automated metrics versus human metrics.
@@ 6 Related Work
Already in the early days of machine translation the need of evaluating across two dimensions was recognized (Pierce & Carroll, 1966; White & O’Connell, 1993). These were also the preferred metrics for many of the NIST MT evaluation in the early 2000s until there was a shift to single HTER scores (Snover et al., 2009; Habash et al., 2011). A similar trend can be found in the WMT evaluations. Whereas in the first editions adequacy and fluency was used (Koehn & Monz, 2006), this approach was quickly dropped in favor of ranking approaches (Callison-Burch et al., 2008) and later a single direct assessment score was given to each segment (Bojar et al., 2016). The use of MQM scores (Kocmi et al., 2022) allows for a distinction into adequacy and fluency evaluation, as we have done in this paper, but these type of evaluation is not available for all conditions and systems. All usual automatic metrics for machine translation evaluation are accuracy metrics as defined in Section 3. To the best of our knowledge there is not much work around intrinsic automatic metrics for the evaluation of the fluency of translations. Vanmassenhove et al. (2021) propose a series of statistical measures and show that they show different characteristics for machine translated and human produced texts. There has been more attention into classifying translations into naturally occuring texts or translationese, which could be considered a measure of naturalness e.g. Kurokawa et al. (2009); Lembersky et al. (2012); Freitag et al. (2022b).
@@
As we have noted on several occasions, our work is heavily inspired by the work of Blau & Michaeli (2018). Their original work has since been significantly extended and analysed in the context of data compression, see Blau & Michaeli (2019); Theis & Wagner (2021); Chen et al. (2022). We extend the framework by considering the translation setting and allowing to choose the reference distribution $R_{\mathbf{y}}$. One important aspect we left mostly unexplored, but which is an important next step once the tradeoff is recognised, is optimising it directly. In the image compression community, this has been recognised for a while, with many state-of-the-art compression models trained using a mixture of accuracy and naturalness (via adversarial losses), see e.g., HiFiC (Mentzer et al., 2020). However, adversarial approaches have recently been far less prominent in the text generation space, due to the difficulty in back-propagating through the generated text. Indeed, finding even the right framework for optimising no-reference metrics is an open problem (Theis, 2024).

### v3

@@ 3.1 The Tradeoff
To make the inevitability of the tradeoff more concrete, consider translating the German sentence “Meine Lehrerin ist nett” to English. A sensible translation would be “My teacher is nice,” which sounds natural, but loses the information from the German sentence that the teacher is female. An alternative translation could be “My female teacher is nice,” which is now fully accurate, but sounds quite unnatural in English. Another, naturally occurring example of this phenomenon is the translation of Japanese honorifics. A completely accurate translation into English would need to introduce diverse paraphrases denoting the social relation between the speakers, which would result in unnatural text. This type of strict semantic differentiation is currently not explicitly assessed by any metric that we are aware of, but might be relevant in certain scenarios.
@@ 4 Translation evaluation in practice
Before moving to the empirical analysis, we lay out the basis for a practical evaluation pipeline. Thus, let ${\mathbf{X}}=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\ldots,{\mathbf{x}}_{N}\}$ and ${\mathbf{Y}}^{ref}=\{{\mathbf{y}}_{1},{\mathbf{y}}_{2},\ldots,{\mathbf{y}}_{N}\}$ be a parallel test set, with source-reference pairs drawn iid: ${\mathbf{x}}_{i},{\mathbf{y}}_{i}\sim P_{{\mathbf{x}},{\mathbf{y}}}$. Furthermore, let $R_{\mathbf{y}}$ be a monolingual reference distribution in the target language, which we could obtain by estimating it from some monolingual corpus.Finally, we consider a set of $M$ translation systems $\mathcal{Q}=\{Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{1},\ldots Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{M}\}$ to be evaluated. Following standard practice, we wish to compare models based on their outputs. To this end, we require that each system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{m}\in\mathcal{Q}$ translates the source texts in the test set and produces a translation set ${\mathbf{Y}}^{m}=\{{\mathbf{y}}_{n}\sim Q_{{\mathbf{y}}\mid{\mathbf{x}}_{n}}^{m}\mid{\mathbf{x}}_{n}\in{\mathbf{X}}\}$. Our goal then is to estimate the accuracy and naturalness of each system using $\mathbf{X}$}, ${\mathbf{Y}}^{ref}$, $R_{\mathbf{y}}$ and the ${\mathbf{Y}}^{m}$s.
@@ 4.1 No-reference metrics as statistical distances
So far, we followed a bottom-up approach, and developed our theory of accuracy and naturalness from the ground up. In this section, we switch to a top-down view, and ask whether current evaluation practices for assessing naturalness conform to our theory. The prevailing practical approach for data quality assessment are no-reference metrics. These can be either human-generated, such as the fluency scores derived from MQM assessments for text [^3] Strictly speaking, MQM evaluation of fluency is not completely reference-free, as the judges have access to the source sentence. But fluency-related scores are judged mainly on a monolingual level. (Freitag et al., 2021), or automated, such as the perplexity of a language model for text. We formalise no-reference metrics by assuming we have a set of $K$ critics $\mathcal{C}=\{f_{1},f_{2},\ldots f_{K}\}$ (usually with $K=1$), each of which assigns a score to each set of translations ${\mathbf{Y}}^{m}$. Then, for each $f_{k}\in\mathcal{C}$, we compute the score $s_{k}^{m}=N^{-1}\sum_{{\mathbf{y}}\in{\mathbf{Y}}^{m}}^{N}f_{k}({\mathbf{y}})$ and average the result to get the final score $s^{m}=K^{-1}\sum_{k=1}^{K}s_{k}^{m}$. Within the context of MQM assessment, the critics could be the set of human raters evaluating the output of a translation system. Then, $f_{k}({\mathbf{y}})$ could be the $k$th rater’s fluency score for the hypothesis $\mathbf{y}$} in ${\mathbf{Y}}^{m}$ derived from their MQM evaluation according to the conversion scheme of Freitag et al. (2021). Does this evaluation procedure fit our definition on naturalness? If so, how? As an answer to this question we propose the following statistical distance between two distributions $Q$ and $R$:
$D_{1}(Q,R\mid\mathcal{P})=\mathbb{E}_{f\sim\mathcal{P}}\left[\left\lvert\mathbb{E}_{Y\sim Q}[f(Y)]-\mathbb{E}_{Y\sim R}[f(Y)]\right\rvert\right$] (4)
where $\mathcal{P}$ is a probability distribution over critics (technically, it is a stochastic process). Note the similarity of Equation˜4 to integral probability metrics (Müller, 1997): instead of maximising, we average over critics instead; see Appendix˜A for a detailed discussion. Formally defining $D_{1}$ is somewhat technical, hence we relegate the details of the definition and the analysis of basic properties to Appendix˜E. Here we argue that averaging no-reference scores is, in fact, a rough Monte Carlo estimate of an appropriately chosen $D_{1}$. Indeed, assuming that the critic scores are all non-negative and are such that the reference distribution $R_{\mathbf{y}}$ always achieves a perfect score for any $f\sim\mathcal{P}$, that is, $\mathbb{E}_{Y\sim R_{\mathbf{y}}}[f(Y)]=0$, then for a test set ${\mathbf{Y}}^{m}$ we have $D_{1}(Q^{m}_{\mathbf{y}},R_{\mathbf{y}}\mid\mathcal{P})\approx(KN)^{-1}\sum_{k=1}^{K}\sum_{{\mathbf{y}}\in{\mathbf{Y}}^{m}}^{N}f_{k}({\mathbf{y}})$ with $f_{k}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{P}$, which is precisely the score $s$ we obtain from averaging no-reference metrics.
@@ 5.1 Approximating the Accuracy-Naturalness curve using LLMs
First, we aim to get a basic feel for the accuracy-naturalness curve. Unfortunately, performing the optimisiation in Equation˜3 is hopeless. Hence, we approximate it in the following way: using Gemini 1.5 Flash (Gemini Team et al., 2024), we generated $K=1024$ candidate translations $\{{\mathbf{y}}_{i}^{k}\}_{k=1}^{K}$ for each source sentence ${\mathbf{x}}_{i}$ in the WMT24 en $\to$ de test set; see Appendix˜F for the precise details. Next, we chose to two accuracy metrics $\Delta$ to test for: chrF (Popović, 2015) and Comet (Rei et al., 2020). For the naturalness metric, we used the pretrained Gemma2 2B model log-perplexities (Gemma Team et al., 2024). Concretely, let $\Gamma(y)$ denote the probability of a sequence of tokens $y$ as predicted by the Gemma2 model, let $\lvert y\rvert$ denote the number of tokens in $y$ and hence let $\mathrm{LPP}(Q)=\mathbb{E}_{Y\sim Q}\left[-\log\Gamma(Y)/\lvert Y\rvert\right$] denote the log-perplexity of a distribution $Q$ with respect to $\Gamma$. Then, our concrete instantiation of $D_{1}$ is $D(Q,R)=\left\lvert\mathrm{LPP}(Q)-\mathrm{LPP}(R)\right\rvert$. As we show in Section˜E.1.2, this quantity can also be viewed as a more general version of the $D_{1}$ distance. We estimated the monolingual target reference distribution $R_{\mathbf{y}}$, using 20,000 entries of the 2024 NewsCrawl DEU Monolingual dataset [^5] Available at https://data.statmt.org/news-crawl/de/.. Of course this choice for $R_{\mathbf{y}}$ represents only a specific type of language (that of news text), and does not represent "colloquial German" for example. Depending on the application, $R_{\mathbf{y}}$ could be adapted by using a different dataset; in our case it merely serves an illustrative purposes In our experiments, we found that $\mathrm{LPP}(R_{\mathbf{y}})$ was always smaller than $\mathrm{LPP}(Q^{m}_{\mathbf{y}})$, for each system we examined, hence $D$ is equivalent up to an additive constant to the average model log-perplexity on the test dataset $D(Q_{\mathbf{y}},R_{\mathbf{y}})=\mathrm{LPP}(Q_{\mathbf{y}})+\mathcal{O}(1)$. To optimise Equation˜3, we first construct the objective’s “one-shot Lagrange dual” with tradeoff parameter $\beta$:
@@
We plot the results in Figure˜1 and see that it verifies our theory: 1) there is indeed a tradeoff curve (see also Theorem˜B.1) and 2) the curve is non-increasing and concave, as ˜3.1 predicts. Finally, note that curves produced this way overestimate the true AN function and hence are overly optimistic, see Appendix˜F for the details.
@@ 5.2 Evaluating existing translation systems using the accuracy-naturalness framework
The plots of the previous section depict a simulation of the tradeoff curve through a controlled experiment, but how do real systems behave with respect to the ideal curve? In the top row of Figure˜2, we include points for the systems that participated in the shared task in addition to the previously computed curve. All systems lie below the curve, as expected. Furthermore, the best performing systems in the evaluation (according to human judgement) are the ones that are closer to the curve. In order to investigate this assertion in more detail, in the bottom row of Figure˜2 and in Figure˜3 we compare human fluency scores against automated accuracy metrics and human adequacy scores [^6] See Appendix D for details on the computation of fluency and adequacy scores from MQM evaluations. for those languages for which MQM evaluations are available. Note that in this case we cannot compute (or simulate) the tradeoff curve for human judgements.
@@
Figure 2: System performances on the WMT24 en $\to$ de test set. $\bullet$ marks represent LLMs, $\blacksquare$ marks represent systems specifically trained for MT, $\blacktriangle$ marks represent “online” systems and the $\blacklozenge$ represents the performance of an alternative human reference. For further explanation of these categories, please see Kocmi et al. (2024). Top row: “Fully automatic” accuracy-naturalness comparisons using chrF and Comet as accuracy metrics and Gemma2B log-perplexity (see section˜5.1) as the naturalness metric. We also include the approximate AN curves from Figure˜1, though note that it is an optimistic estimate of the real curve. Bottom row: comparisons of automatic accuracy metrics (chrF and Comet) against human MQM fluency ratings, as described in section˜5.2.
Figure 3: Plot of the accuracy-naturalness tradeoff derived from human MQM adequacy and fluency ratings, as described in Section˜5.2. We highlight the top-performing systems at the Pareto frontier.
@@
Secondly, we focus on interpreting the results for the performance of single systems. Unbabel (Rei et al., 2024) had the dominant system in the automatic metrics, but in human evaluations other systems were judged better. Unbabel’s system was clearly optimized for Comet, an accuracy measure. Looking at the plots in Figure˜3 we can see that the Unbabel system has a quite strong adequacy score, but due to the tradeoff the fluency of the system suffers. Other systems that human raters judged to be better lie closer to the tradeoff curve.
@@ 6 Related Work
Already in the early days of machine translation the need of evaluating across two dimensions was recognized (Pierce & Carroll, 1966; White & O’Connell, 1993). These were also the preferred metrics for many of the NIST MT evaluation in the early 2000s until there was a shift to single HTER scores (Snover et al., 2009; Habash et al., 2011). A similar trend can be found in the WMT evaluations. Whereas in the first editions adequacy and fluency was used (Koehn & Monz, 2006), this approach was quickly dropped in favor of ranking approaches (Callison-Burch et al., 2008) and later a single direct assessment score was given to each segment (Bojar et al., 2016). The use of MQM scores (Kocmi et al., 2022) allows for a distinction into adequacy and fluency evaluation, as we have done in this paper, but these type of evaluation is not available for all conditions and systems. All usual automatic metrics for machine translation evaluation are accuracy metrics as defined in Section˜3. To the best of our knowledge there is not much work around intrinsic automatic metrics for the evaluation of the fluency of translations. Vanmassenhove et al. (2021) propose a series of statistical measures and show that they show different characteristics for machine translated and human produced texts, but these do not constitute statistical distances as we require here. There has been more attention into classifying translations into naturally occuring texts or translationese, which could be considered a measure of naturalness e.g. (Kurokawa et al., 2009; Lembersky et al., 2012; Freitag et al., 2022b).
@@
As we have noted on several occasions, our work is heavily inspired by the work of Blau & Michaeli (2018). Before its publication, the neural image restoration and compression community was grappling with the analogous issue and widely held the same misconception the MT community holds now: despite developing ever-more sophisticated, so called “perceptual distortion metrics” that were inspired by/based on the human visual system, and which correlated well with human judgments, systems optimized using these metrics inevitably showed undesirable visual artefacts. The community’s response to this was that “we just haven’t found the right metric yet,” and spent considerable effort on solving this issue. Blau and Michaeli’s work showed that this effort is hopeless, and since then evaluating the performance of image reconstruction and compression algorithms along the distortion-perception plane has become standard practice. We hope our work will help bring about a similar change in the MT community.
The original idea has since been significantly extended and analysed in the context of data compression, see Blau & Michaeli (2019); Theis & Wagner (2021); Chen et al. (2022). We extend the framework by considering the translation setting and allowing to choose the reference distribution $R_{\mathbf{y}}$. One important aspect we left mostly unexplored, but which is an important next step once the tradeoff is recognised, is optimising it directly. In the image compression community, this has been recognised for a while, with many state-of-the-art compression models trained using a mixture of accuracy and naturalness (via adversarial losses), see e.g., HiFiC (Mentzer et al., 2020). However, adversarial approaches have recently been far less prominent in the text generation space, due to the difficulty in back-propagating through the generated text. Indeed, finding even the right framework for optimising no-reference metrics is an open problem (Theis, 2024). In the field of machine translation and text generation there has been some work on optimizing several objectives concurrently, e.g. (Duh et al., 2012; Kumar et al., 2021). Similar frameworks can potentially be used to optimize for a specific tradeoff of adequacy and fluency.